Williams Optical Inc. is considering a new lean product cell. The present manufacturing approach produces a product in four separate steps. The production batch sizes are 61 units. The process time for each step is as follows:

Process Step 1 6 minutes
Process Step 2 8 minutes
Process Step 3 5 minutes
Process Step 4 4 minutes

The time required to move each batch between steps is 15 minutes. In addition, the time to move raw materials to Process Step 1 is also 15 minutes, and the time to move completed units from Process Step 4 to finished goods inventory is 15 minutes.

The new lean layout will allow the company to reduce the batch sizes from 61 units to 4 units. The time required to move each batch between steps and the inventory locations will be reduced to 3 minutes. The processing time in each step will stay the same.

Determine the value-added, non-value-added, and total lead times, and the value-added ratio under the present and proposed production approaches. If required, round percentages to one decimal place.

Present Approach Proposed Approach
Value-added time min min
Non-value-added time min min
Total lead time min min
Value-added ratio (as a percent) % %
Answer #1

ANSWER:

As per the given Question

Present Approach

Value added time = Process time Step 1 + Process time Step 2 + Process time Step 3 + Process time Step 4

= 6 + 9 + 4 + 8 = 27 minutes

Non-value added time = Total within batch wait time + Move time

= [Total time required for four steps *(Batch size -1)] + Move time

= [27 * (70-1)] + 60 mins

= 1,923 minutes

Total lead time = Value added time + Non-value added time

= 27 + 1,923 = 1,950 minutes

Value added ratio = Value added time / Total lead time

= 27 / 1,950 = 1.4%

Proposed Approach

Value added time = Process time Step 1 + Process time Step 2 + Process time Step 3 + Process time Step 4

= 6 + 9 + 4 + 8 = 27 minutes

Non-value added time = Total within batch wait time + Move time

= [Total time required for four steps *(Batch size -1)] + Move time

= [27 * (4-1)] + 10 mins

= 91 mins

Total lead time = Value added time + Non-value added time

= 27 + 91 = 118 mins

Value added ratio = Value added time / Total lead time

= 27 / 118 = 22.9%

Thank you for given Question......kindly rate.......it helps me a lot